OpenStudy (anonymous):

Choose the best definition for the following term: substitution. 1) A letter that holds the place for some unknown value in mathematics 2) A process where you must look for terms that have identical variable parts and then combine their coefficients 3) A process where if two things are equal, one can be put in the place of the other and nothing will change 4) Terms that have identical variable parts Save

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi Misty!

OpenStudy (misty1212):

i would go with C since a) it is almost always C and b) it is the only answer that makes sense

OpenStudy (anonymous):

I thougth substitution meant the act of replacing the letters with actual numbers. Like subbing values in

OpenStudy (anonymous):

However C is DEFINITELY CORRECT and I CAN SWEAR to my mother.

OpenStudy (misty1212):

you can substitute variables for sure \[a^2-b^2=(a+b)(a-b)\] put \(a=2x,b=y^2\) get \[4x^2-y^4=(2x+y^2)(2x-y^2)\]

OpenStudy (anonymous):

How can you type up equations so fast?

OpenStudy (misty1212):

\[\LaTeX\]

OpenStudy (anonymous):

obrigadooooo

OpenStudy (misty1212):

de nada \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

hablas espaniol? Muy bien yo tambien